# An Innovative Attention-based Ensemble System for Credit Card Fraud Detection

### v1

@@ III-D CCF detection Framework: Structure, Development, and Examination
This study enhances the traditional stacking architecture by incorporating an attention layer and a selection layer. Fig. 5 shows the block diagram of the proposed architecture for CCF detection. In the following, we delve into the details of this entirely attention-based parallel architecture.
III-D1 First layer classifier selection
Selecting first layer classifiers is a crucial step in building an effective stacking model. The diversity and performance of these classifiers have a significant impact on the stacking model’s overall performance. Here are key considerations and strategies for selecting first layer classifiers (base models) in a stacking classifier:
Performance
Using feeble models can harm the overall performance, as they may introduce too much noise. However, models do not need to be top performers individually, as their errors might be corrected by the meta-learner. The best candidates for first layer classifiers are those not only with the best performance individually, but also whose predictions possess an acceptable level of diversity. Under these circumstances, the classifiers will complement each other well. Table I shows the 6 base classifiers that were able to meet the above criteria.
Diversity
Diverse base classifiers are crucial for making errors in different parts of the dataset. Similar classifiers lead to similar mistakes, reducing stacking effectiveness. Correlation between the predictions of first layer classifiers is a measure for assessing the level of diversity. Fig. 6 shows the correlation values between the predictions for each pair of first-level classifiers. It highlights how these two classifiers’ predictions are connected or related. This could be visualized using lines, colors, or patterns indicating areas, with medium or low correlation. High correlation areas indicate that both classifiers typically agree on sample classification, while low correlation areas suggest disagreement.
Figure 6: The correlation among the predictions generated by the classifiers in the first layer.
III-D2 attention layer
At the heart of the suggested design lies the attention layer, which allows the model to concentrate on classifiers. This layer allows the model to assign separate weightings to the predictions of each classifier based on the belief associated with each classifier. Considering the DOWA aggregation operator, which assigns a negligible weight to the classifier prediction that is an outlier compared to the other two classifiers, it is necessary for the classifiers, whose predictions are to be fused by this operator, to have sufficiently close correlations in their predictions. For example, if we assume the correlation between classifiers $C1$, $C2$, and $C3$ to be $C1C2:0.85$, $C1C3:0.9$, and $C2C3:0.4$, then by aggregating the predictions of these three classifiers using the DOWA operator, this operator will disregard the impact of one of the two classifiers, either $C2$ or $C3$, for each sample. Therefore, to employ this operator in its best performance, we select the classifiers Boosted Trees (BT), Extra Trees (ET), and AdaBoost (AB), which according to Fig. 6 have prediction correlations of 0.69, 0.64, and 0.56, respectively, which are significantly close to each other, for aggregation using DOWA. The predictions of the remaining three classifiers, namely SGD, SVM, and MLP, are aggregated using the IOWA operator.
Therefore, the weights are calculated for the three classifiers BT, ET, and AB by the DOWA aggregation operator, and for the three classifiers SGD, SVM, and MLP by the IOWA operator. Ultimately, the fusion of the probabilities assigned to each sample by the first layer classifiers is done using the calculated weights, and these probabilities are sorted in descending order. The output of the DOWA operator for each sample $i=1,2,...,N$ is a vector in the form of
$F_{DOWA}=[f(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0}),f(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})]^{T},$ (17)
and similarly, the output of the IOWA operator is a vector in the form of
$\small F_{IOWA}=[f(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0}),f(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})]^{T},$ (18)
where $p_{BT}^{0}$indicates the probability of sample i belonging to class 0 as classified by BT, and $f(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})$ represents the fused value of absolute probabilities assigned to class zero by the three classifiers BT, ET, and AB. Considering that the DOWA and IOWA aggregation operators each aggregate the soft outputs of the three specified classifiers, it can be said that each of these operators is a mapping from the space $R^{6}$ to $R^{2}$.
III-D3 selection layer
This layer receives two vectors, $F_{DOWA}$ and $F_{IOWA}$, as inputs. The basis of its operation for each sample is according to the Algorithm 1.
Algorithm 1 : Sample-Based Aggregated Value Selector
Input: $f_{i}(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})$, $f_{i}(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})$, $f_{i}(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0})$, $f_{i}(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})$
@@ III-D1 First layer classifier selection
In fact, for each sample, this layer decides to intelligently select one of the two vectors, $F_{DOWA}$ or $F_{IOWA}$. The logic of selection is that whichever of the two operators, DOWA or IOWA, has greater confidence in a sample belonging to one of the two classes, 0 or 1, the aggregated value by that operator is chosen. In this way, we use a smart selection of aggregated values by both operators to feed the meta-learner. It can be said that the selection layer synergistically combines the results from these two operators.
III-D4 Meta-learner selection
The stacking process is significantly influenced by a meta-learner, as it plays a crucial role in determining the classifier’s performance. The selection of a classifier for meta learning is crucial as it directly influences the stacked ensemble’s generalization on unseen instances and, subsequently, its overall predictive accuracy. In the context of the stacking architecture, the Ridge classifier can be considered a favorable option due to its moderate complexity, which facilitates interpretability. Additionally, it incorporates regularization techniques to mitigate overfitting and effectively handles potential multicollinearity among the features used for learning. In Table II, the test results are presented, indicating that the Ridge classifier outperforms other base classifiers.
TABLE I: The degree of accuracy achieved by the first layer classifiers when evaluated on the test data.
Classifier Accuracy
Boosted Trees (BT) $93.64\pm 0.22\%$
Stochastic Gradient Descent (SGD) $92.48\pm 0.13\%$
Extra Trees (ET) $94.09\pm 0.31\%$
AdaBoost (AB) $95.33\pm 0.19\%$
Support vector Machine (SVM) $94.77\pm 0.25\%$
Multi-Layer Perceptron (MLP) $95.87\pm 0.23\%$
TABLE II: Accuracy metric for choosing the meta-learner.
Learner Accuracy
Decision Tree $95.32\pm 0.16\%$
Gradient Boosting $96.71\pm 0.13\%$
K-nearest Neighbor $96.83\pm 0.11\%$
Logistic Regression $98.44\pm 0.19\%$
Ridge Classifier $99.95\pm 0.04\%$
CatBoost $97.63\pm 0.19\%$
IV Findings and Analysis
In this segment, we delve deeper into the results and present a range of performance indicators to evaluate the effectiveness of the proposed algorithm. These include measures such as accuracy, sensitivity (also known as recall), precision, specificity, Matthew’s Correlation Coefficient (MCC), and the F1-Score.
IV-A Performance metrics
IV-A1 Precision
Precision quantifies the ratio of correct positive predictions to the total number of positive predictions made by a model, highlighting the accuracy of its positive predictions.
$Precision=\frac{TP}{TP+FP}.$ (19)
IV-A2 Specificity
Specificity measures the ratio of accurate negative predictions to the total negative predictions made by a model, underlining the precision of its negative predictions, e.g.,
$Specificity=\frac{TN}{TN+FP}.$ (20)
IV-A3 Accuracy
Accuracy assesses the total proportion of both true positive and true negative predictions made by a model compared to the overall number of predictions, reflecting the general reliability of the model’s predictions, e.g.,
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}.$ (21)
IV-A4 Sensitivity
Sensitivity quantifies the percentage of true positive instances accurately identified by a model, emphasizing the model’s capacity to detect all positive cases, e.g.,
$Sensitivity=\frac{TP}{TP+FN}.$ (22)
IV-A5 MCC
Unlike other metrics that may only focus on the positive class or are overly optimistic in imbalanced datasets, MCC provides a balanced measure, even if the classes are of very different sizes. MCC is defined mathematically as:
$\small MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}.$ (23)
IV-A6 F1-Score
The F1-score is a harmonized measure that considers precision and recall. It offers a singular metric that equilibrates the balance between precision and recall, proving valuable in situations involving datasets with imbalances, e.g.,
$F1-Score=\frac{2\times Precision\times sensitivity}{Precision+sensitivity},$ (24)
where $TP$ is true positive, $FP$ is false positive, $TN$ is true negative and $FN$ is false negative.
To prevent data leakage at all stages, including predictions by first layer classifiers, getting weights through a learning process by the IOWA operator, and the final prediction by the meta-learner, the 10-fold cross-validation method was used. The exceptional performance of the proposed architecture on the test data aligns with the findings reported in Table III.
Fig. 7 presents the receiver operating characteristic (ROC) curve, which showcases the correlation between the true positive rate (TPR) and the false positive rate (FPR) across different classification thresholds. The ROC curves for both classes demonstrate a significant level of overlap, suggesting comparable performance between the two classes. The ROC curve’s area under the curve (AUC) functions as a metric to assess classifier performance. A higher AUC value, approaching 1, signifies superior classifier performance. The suggested model demonstrates the AUC value of 1.
Figure 7: ROC curve along with corresponding AUC on test data.
TABLE III: Performance metrics of the proposed method.
Metric Value
Precision $100\%$
Accuracy $99.95\%$
Sensitivity $99.90\%$
Specificity $100\%$
MCC $99.91\%$
F1-Score $99.94\%$
V Prospects for Future Studies

### v2

@@ II-D CCF detection Framework: Structure, Development, and Examination
This study enhances the traditional stacking architecture by incorporating an attention layer and a confidence-aware combination layer. Fig. 3 shows the block diagram of the proposed architecture for CCF detection. In the following, we delve into the details of this entirely attention-based ensemble architecture.
II-D1 Diversity-Based Selection Strategy
To evaluate the performance of deep learning models in fraud detection, we compare four architectures: CNN, RNN, LSTM, and GNN. These models are trained and optimized using cross-validation to ensure fair comparison. The main hyperparameters for each method are as follows.
(1) CNN [27]: Number of convolutional layers = 3, Kernel size = 3, Stride = 1, Pooling type = MaxPooling, Pool size = 2, Activation function = ReLU, Fully connected layers = 2 (with 128 and 64 neurons), Optimizer = Adam (learning rate = 0.001), Epochs = 200.
(2) RNN [28]: Number of recurrent layers = 2, Hidden units per layer = 128, Activation function = Tanh, Dropout rate = 0.2, Optimizer = RMSprop (learning rate = 0.001), Epochs = 200,
(3) LSTM [28]: Number of LSTM layers = 2, Hidden units per layer = 128, Activation function = Tanh, Dropout rate = 0.3, Recurrent dropout = 0.2, Optimizer = Adam (learning rate = 0.0005), Epochs = 200.
(4) GNN [29]: Number of graph convolution layers = 3, Hidden units per layer = 64, Graph aggregation method = Mean pooling, Activation function = LeakyReLU, Dropout rate = 0.2, Optimizer = Adam (learning rate = 0.001), Epochs = 200.
Based on the correlation matrix in Fig. 4, we observe that RNN and LSTM have the highest correlation (0.87), indicating considerable overlap in their predictive patterns. Consequently, combining them would offer limited diversity and might not significantly improve overall performance. By contrast, CNN and RNN (correlation: 0.71) exhibit moderate similarity, providing enough redundancy for reliability while maintaining enough difference for enhanced coverage. Similarly, GNN and LSTM show the lowest correlation (0.55), suggesting that each model captures distinct aspects of the data and could thus complement one another well. Therefore, in our framework, CNN and RNN are aggregated via the DOWA operator, while GNN and LSTM are combined using the IOWA operator. This diversity-based selection ensures a balanced trade-off between redundancy (stability) and complementarity (robustness), ultimately improving the ensemble’s ability to detect fraud across a variety of transaction patterns.
Figure 4: Correlation matrix illustrating the diversity among predictions from RNN, GNN, CNN, and LSTM. Darker shades indicate higher agreement, whereas lighter shades reflect lower correlation.
II-D2 Attention Layer: DOWA and IOWA Combinations
In our proposed framework, we first introduce an attention layer that uses the DOWA operator to combine the soft outputs of CNN and RNN. DOWA assigns higher weights to classifiers whose predictions lie closer to the overall mean, preventing any single classifier with inconsistent outputs from excessively skewing the final aggregate. Since CNN and RNN exhibit a moderate correlation (0.71), this approach balances reliability and diversity by emphasizing their areas of agreement while retaining their distinct predictive signals, ultimately yielding a more robust and stable ensemble.
Building on this, we extend the attention layer by incorporating the IOWA operator to merge the soft outputs of LSTM and GNN. In contrast to DOWA, which highlights predictions near the mean, IOWA weights inputs according to their order and assigned significance, making it especially suitable for combining classifiers with low correlation. Given that LSTM and GNN demonstrate the lowest correlation (0.55), this weighted strategy exploits their complementary strengths, LSTM’s ability to capture temporal patterns and GNN’s capacity for graph-structured data representation. By carefully assigning weights through IOWA, our framework amplifies reliable signals from each model while minimizing the influence of outlier predictions, further strengthening the ensemble’s overall performance in fraud detection.
Therefore, the DOWA operator is used to compute weights for the RNN and CNN classifiers, and the IOWA operator does the same for the GNN and LSTM classifiers. Once these weights are learned, the soft probabilities from each first-layer classifier are combined accordingly and sorted in descending order. For a sample $i\in\{1,2,\dots,N\}$, the DOWA operator produces a two-dimensional output of the form $F_{\mathrm{DOWA}}=\bigl{[}f\bigl{(}p_{\mathrm{RNN}}^{0},p_{\mathrm{CNN}}^{0}\bigr{)},\,f\bigl{(}p_{\mathrm{RNN}}^{1},p_{\mathrm{CNN}}^{1}\bigr{)}\bigr{]}^{T},$ while the IOWA operator yields
$F_{\mathrm{IOWA}}=\bigl{[}f\bigl{(}p_{\mathrm{GNN}}^{0},p_{\mathrm{LSTM}}^{0}\bigr{)},\,f\bigl{(}p_{\mathrm{GNN}}^{1},p_{\mathrm{LSTM}}^{1}\bigr{)}\bigr{]}^{T}$.
Here, $p_{\mathrm{RNN}}^{0}$ denotes the probability that sample $i$ belongs to class 0 as predicted by the RNN, and $f\bigl{(}p_{\mathrm{RNN}}^{0},p_{\mathrm{CNN}}^{0}\bigr{)}$ represents the fused value of the respective class-0 probabilities from the RNN and CNN. Because each of these operators aggregates the outputs of two classifiers, they effectively map from $\mathbb{R}^{4}$ (the four class probabilities, two from each classifier) to a reduced space $\mathbb{R}^{2}$.
II-D3 Confidence-Aware Combination (CAC) Layer
After generating two fused probability vectors, $F_{\mathrm{DOWA}}$ from the CNN–RNN pair and $F_{\mathrm{IOWA}}$ from the GNN–LSTM pair, the framework applies a confidence-based mechanism to decide which set of fused probabilities is passed to the meta-learner. Specifically, we quantify uncertainty by comparing the absolute difference between class-0 and class-1 probabilities for each aggregator. Formally, for each sample $i$, $\Delta_{\mathrm{DOWA}}=\Bigl{\lvert}f\bigl{(}p_{\mathrm{RNN}}^{0},\,p_{\mathrm{CNN}}^{0}\bigr{)}\;-\;f\bigl{(}p_{\mathrm{RNN}}^{1},\,p_{\mathrm{CNN}}^{1}\bigr{)}\Bigr{\rvert}\quad\text{and}\quad\Delta_{\mathrm{IOWA}}=\Bigl{\lvert}f\bigl{(}p_{\mathrm{GNN}}^{0},\,p_{\mathrm{LSTM}}^{0}\bigr{)}\;-\;f\bigl{(}p_{\mathrm{GNN}}^{1},\,p_{\mathrm{LSTM}}^{1}\bigr{)}\Bigr{\rvert}.$
We interpret the operator with the smaller absolute difference as having higher uncertainty. Consequently, if $\Delta_{\mathrm{IOWA}}\ \leq\Delta_{\mathrm{DOWA}}$, we feed $F_{\mathrm{DOWA}}$ (the DOWA-fused output) into a multi-layer perceptron (MLP) meta-learner; otherwise, we select $F_{\mathrm{IOWA}}$. As detailed in Algorithm 1, this decision process is repeated for each sample. By always routing the more confident fused probabilities to the meta-learner, the final classification performance is enhanced, particularly under challenging conditions such as imbalanced datasets.
Algorithm 1 : Confidence-Aware aggregator selection choosing between the DOWA- and IOWA-fused outputs based on uncertainty measures for each sample.
Input: $f_{i}(p_{RNN}^{0},p_{CNN}^{0})$, $f_{i}(p_{RNN}^{1},p_{CNN}^{1})$, $f_{i}(p_{GNN}^{0},p_{LSTM}^{0})$, $f_{i}(p_{GNN}^{1},p_{LSTM}^{1})$
@@ II-D3 Confidence-Aware Combination (CAC) Layer
II-D4 Meta-Learner Configuration
In our final classification stage, we employ a MLP as the meta-learner to process the fused probabilities (i.e., $F_{\mathrm{DOWA}}$ or $F_{\mathrm{IOWA}}$ selected by the confidence-aware combination layer. The MLP architecture consists of two hidden layers, each containing $32$ neurons with ReLU activation, to capture non-linear relationships between the fused features. A dropout rate of $0.2$ is applied to mitigate overfitting and improve generalization. We train the MLP using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of $64$, and early stopping based on a validation loss threshold. This configuration ensures both effective learning and a balance between model complexity and computational efficiency.
Figure 5: Confusion matrix for Dataset 1, using the proposed ensemble model with an 80-20 stratified train-test split.
Figure 6: Performance comparison of CNN, RNN, LSTM, GNN, and the proposed ensemble model on Dataset 1.
Figure 7: Confusion matrix for Dataset 2, using the proposed ensemble model with an 80-20 stratified train-test split.
Figure 8: Performance comparison of CNN, RNN, LSTM, GNN, and the proposed ensemble model on Dataset 2.
Figure 9: Confusion matrix for Dataset 3, where the ensemble model was trained on balanced Dataset 1 and tested on imbalanced Dataset 2.
Figure 10: Performance comparison of CNN, RNN, LSTM, GNN, and the proposed ensemble model on Dataset 3.
Figure 11: Impact of incorporating the attention and confidence-aware combination layers in the stacking model on the test data of Dataset 1.
Figure 12: Impact of incorporating the attention and confidence-aware combination layers in the stacking model on the test data of Dataset 2.
III Results
In this section, we delve deeper into the results of our proposed algorithm and present a comprehensive set of performance metrics, including $A_{c}$, recall ($R_{e}$), precision ($P_{r}$), specificity ($S_{p}$), and the F1-Score ($F_{1}$). The subsequent figures illustrate how the attention-based ensemble model performs on three different datasets. For each dataset, we provide two visualizations: a confusion matrix, which displays classification outcomes for the test set, green areas indicate correct predictions and red areas denote misclassifications, and a performance comparison graph, contrasting the ensemble’s key evaluation metrics against those of the individual deep learning models (CNN, RNN, LSTM, and GNN). The fitted probability threshold is shown in each confusion matrix, defining the boundary for classifying instances as either fraudulent (1) or legitimate (0). Additionally, a bar chart adjacent to each matrix offers a quick snapshot of how many samples fall into each category. This side-by-side comparison underscores the benefits of our attention-based approach in terms of $A_{c}$, $R_{e}$, $P_{r}$, $S_{p}$, and the $F_{1}$.
Dataset 1 (balanced): In Fig. 5 and Fig. 6, the model is trained and tested with an 80–20 stratified split on a balanced dataset (Dataset 1). As shown in the confusion matrix (Fig. 5), the proposed model achieves a high true negative rate, effectively identifying legitimate transactions. However, the number of correctly identified fraud cases remains limited, an expected result given the overall lower prevalence of fraud in this dataset. The performance comparison (Fig. 6) demonstrates that the ensemble model consistently surpasses individual deep learning models, with a particularly strong showing in $R_{e}$, a critical metric for fraud detection.
Dataset 2 (heavily imbalanced): For Dataset 2 (Fig. 7 and Fig. 8), the same 80–20 stratified split is applied, but the task becomes more challenging due to a significantly skewed class distribution. The confusion matrix (Fig. 7) reveals an increased number of false negatives, indicating that detecting fraudulent transactions is harder under these conditions. This likely stems from distinct fraud patterns and their lower representation in Dataset 2. Nonetheless, the ensemble model still outperforms each individual model, as evident in Fig. 8, maintaining higher $R_{e}$ and $S_{p}$, even under difficult, highly imbalanced circumstances.
Dataset 3 (cross-dataset evaluation): In Dataset 3 (Fig. 9 and Fig. 10), we adopt a different training strategy: Dataset 1 (balanced) is used for training, while Dataset 2 (imbalanced) is reserved for testing. The confusion matrix (Fig. 9) shows that, although the model tends to be conservative, leading to a higher false positive rate, it nonetheless captures a substantial number of fraudulent transactions. Fig. 10 confirms that, despite this trade-off, our ensemble model continues to outperform the standalone classifiers, demonstrating a strong balance between $R_{e}$ and $S_{p}$ when transitioning from balanced training data to highly imbalanced real-world scenarios.
Fig. 11 and Fig. 12 evaluate the impact of incorporating attention-based fusion and confidence-aware combination layers into the stacking model by comparing two versions: a baseline model without these layers and an enhanced version that integrates them. The results show that adding these extra layers significantly improves performance. This improvement indicates that intelligently combining model outputs and reducing uncertainty in final predictions enhances fraud detection effectiveness.
TABLE II: Performance comparison of various ensemble learning models and the proposed attention-based stacking model for CCF detection across multiple evaluation metrics.
Datasets LightGBM RF XGBoost DF AdaBoost ET BDT Proposed
Dataset 1 $P_{r}$ 0.9313 0.8727 0.9280 0.9238 0.9281 0.8853 0.8815 0.9984
$R_{e}$ 0.8606 0.9172 0.8172 0.8592 0.8595 0.8931 0.9114 1.0000
$F_{1}$ 0.8945 0.8942 0.8693 0.8901 0.8923 0.8885 0.8959 0.9992
Dataset 2 $P_{r}$ 0.8847 0.8158 0.8814 0.8799 0.8875 0.7433 0.8375 0.9091
$R_{e}$ 0.7596 0.8204 0.7400 0.7298 0.7596 0.8827 0.8327 0.8081
$F_{1}$ 0.8172 0.8181 0.8042 0.7975 0.8162 0.8079 0.8347 0.8556
Dataset 3 $P_{r}$ 0.0003 0.0012 0.0008 0.0006 0.0018 0.0007 0.0010 0.0095
$R_{e}$ 0.7155 0.7850 0.7395 0.7104 0.7157 0.7785 0.7874 0.8069
$F_{1}$ 0.0005 0.0023 0.0015 0.0012 0.0035 0.0014 0.0019 0.0187
Table II presents a comparative evaluation of various ensemble learning models, including LightGBM, Random Forest (RF), XGBoost, Deep Forest (DF), AdaBoost, Extra Trees (ET), and Boosted Decision Trees (BDT), against the proposed attention-based stacking model for CCF detection. The models are assessed across three datasets and multiple evaluation metrics: $P_{r}$, $R_{e}$, and $F_{1}$. The proposed model consistently outperforms others, achieving the highest $P_{r}$, $R_{e}$, and $F_{1}$, especially in Dataset 1 ($F_{1}$ = 0.9992) and Dataset 2 ($F_{1}$ = 0.8556). Since the features in two datasets were anonymized, we assumed their equivalence when training on Dataset 1 and testing on Dataset 2. However, differences in underlying distributions due to potential variations in feature anonymization may have contributed to the lower $F_{1}$ score observed in Dataset 3 ($F_{1}$ = 0.0187). This suggests that while the proposed model generalizes well across datasets with similar feature representations, discrepancies introduced by anonymization could affect its ability to transfer knowledge effectively across different datasets.
Overall, the results across all three datasets confirm the effectiveness and adaptability of the proposed ensemble approach. On the balanced dataset (Dataset 1), the model demonstrates high reliability, particularly in correctly identifying legitimate transactions and maintaining a strong $R_{e}$ rate. When faced with a heavily imbalanced dataset (Dataset 2), the method preserves its comparative advantage over individual models, underscoring its robustness in challenging real-world conditions. Finally, in the cross-dataset evaluation (Dataset 3), the ensemble model again proves its versatility, striking an efficient trade-off between $R_{e}$ and $S_{p}$ when transitioning from a balanced training set to testing on imbalanced data.
IV Discussion
Fraud detection, which has gained significant research interest in recent years, has been tackled using a wide range of methodologies and algorithms. The following paragraphs highlight notable studies in this area.
Fanai et al. [30] leveraged a deep Autoencoder for representation learning. The Autoencoder reduced the dimensionality of the data, extracting compact and informative representations, which were then used to train models, such as DNN, RNN, and a hybrid CNN-RNN. The Bayesian optimization algorithm was employed to fine-tune hyperparameters.
Tian et al. [31] employed adaptive sampling and aggregation-based graph neural network (ASA-GNN). This model addressed challenges such as noisy data, fraudsters’ camouflage behavior, and the over-smoothing issue in GNNs. ASA-GNN employed a neighbor sampling strategy that used cosine similarity to filter irrelevant nodes.
Zhu et al. [32] proposed a noisy-sample-removed under-sampling scheme for imbalanced classification. This approach addressed the negative impact of noisy samples on classifier performance. The method utilized K-means clustering to remove noisy samples from both majority and minority classes.
The study [33] proposed an attention-based ensemble model. The model improved generalization and reduce variance by introducing a feature-wise attention mechanism and feature diversity regularization. The attention mechanism learned the most informative features, while the regularization term optimized the balance between bias and variance.
The study [34] focus on a federated learning approach which enabled the development of a global model by aggregating locally computed updates from distributed datasets without sharing raw data, ensuring privacy. To tackle class imbalance, the study implemented both individual and hybrid resampling techniques, such as synthetic minority over-sampling technique (SMOTE) and adaptive synthetic (ADASYN).
Zhu et al. [35] proposed an adaptive heterogeneous framework integrating deep RL and attention mechanisms. It used a reinforcement reward mechanism to select source domain instances and a CNN with an attention module to extract semantic features.
Tang et al. [36] concentrated on a federated graph learning approach, combining federated learning and GNN to model transaction relationships while preserving privacy. A graph extension algorithm and adaptive aggregation improved cross-institutional collaboration and addressed data imbalance.
TABLE III: Comparison summary of the proposed method with existing approaches.
Work Dataset Objective Deep Learning Model Performance
2022[6]
Public: 284 807 samples,
492 fraudulent cases
Private: 5 120 000 transactions,
140 000 fraudulent records
Binary
Time-aware
historical-attention-
based LSTM
$P_{r}$ = 0.504
$R_{e}$ = 0.996
$F_{1}$ = 0.669
$P_{r}$ = 0.921
$R_{e}$ = 0.972
$F_{1}$ = 0.946
2023[30]
Public: 284 807 samples,
492 fraudulent cases
Private: 1000 samples,
30% labled
as bad credit
Binary
Deep neural network
+ RNN
+ Hybric CNN + RNN
$F_{1}$ = 0.837
AUC = 0.921
$F_{1}$ = 0.807
AUC = 0.908
2023[31]
Private: 5 120 000 samples,
140 000 fraudulent cases
Public: 160 764 samples,
(28% fraudulent cases)
Public: 20 000 samples,
(balanced dataset)
Binary
Aggregation-based
graph neural
network
$R_{e}$ = 0.884
$F_{1}$ = 0.904
AUC = 0.922
$R_{e}$ = 0.737
$F_{1}$ = 0.774
AUC = 0.835
$R_{e}$ = 0.731
$F_{1}$ = 0.593
AUC = 0.571
2023[32]
Private: 614 samples,
192 fraudulent cases
Private: 3 075 samples,
448 fraudulent cases
Private: 30 000 samples,
6 636 fraudulent cases
Binary
Decision tree +
Logistic regression +
Random forest
as base classifier
in the NUS framework
$F_{1}$ = 0.890
G-mean = 0.910
$F_{1}$ = 0.860
G-mean = 0.870
$F_{1}$ = 0.830
G-mean = 0.850
2023[33]
Public: 284 807 samples,
492 fraudulent cases
Private: 3 500 000 samples,
fraud rate 1.59%
Binary
A feature-wise
attention-based
boosting ensemble
model
$F_{1}$ = 0.885
$F_{1}$ = 0.929
2024[34]
Public: 284 807 samples,
492 fraudulent cases
Binary
Federated CNN
models and
base classifiers
$A_{c}$ = 0.999
$P_{r}$ = 0.826
$R_{e}$ = 0.809
2024[35]
Private: 5 125 107 samples,
147 829 fraudulent cases
Public: 284 807 samples,
492 fraudulent cases
Binary
A CNN with
attention
mechanism
$P_{r}$ = 0.921
$F_{1}$ = 0.946
$P_{r}$ = 0.902
$F_{1}$ = 0.880
2024[36]
Public: 284 807 samples,
492 fraudulent cases
590 540 samples
Binary
Federated graph
learning
$P_{r}$ = 0.925
$R_{e}$ = 0.910
$F_{1}$ = 0.918
$A_{c}$ = 0.935
$P_{r}$ = 0.710
$R_{e}$ = 0.615
$F_{1}$ = 0.640
$A_{c}$ = 0.800
2025[37]
Public: 284 807 samples,
492 fraudulent cases
Binary
Sequential detection:
LSTM
Non-Sequential
detection:
LightGBM
$F_{1}$ = 0.850
AUC = 0.870
$F_{1}$ = 0.870
AUC = 0.960
Our work
Public: 568 630 samples,
284 615 fraudulent cases
Public: 284 807 samples,
492 fraudulent cases
Public: 853 437 samples,
284 807 fraudulent cases
Binary
Attention-based
ensemble of CNNs
and GNN
$A_{c}$ = 0.9992
$F_{1}$ = 0.9992
Pre = 0.9984
$R_{e}$ = 1.0
$A_{c}$= 0.9995
$F_{1}$ = 0.8556
Pre = 0.9091
$R_{e}$ = 0.8081
$A_{c}$ = 0.830
$R_{e}$ = 0.8069
$S_{p}$ = 0.8539
The study [37] developed integrated resampling methods and data-dependent classifiers. It combines one-class SVM with SMOTE and random under-sampling to preserve the distribution of fraud instances. The framework outputs are analyzed using LightGBM for non-sequential fraud detection and LSTM for sequential fraud detection.
Table III compares this study’s approach with other recent fraud detection methods that employ deep neural networks. While many existing solutions have achieved respectable performance, our ensemble framework distinguishes itself through its uncertainty-aware design, combining DOWA, IOWA, and a confidence-driven gating mechanism, to effectively handle both balanced and highly imbalanced datasets. This architecture not only delivers strong $A_{c}$ and $R_{e}$ but also remains computationally feasible, making it adaptable for deployment in real-time payment gateways. Moreover, by integrating SHAP-based feature selection, each chosen attribute can be mapped to concrete transactional insights, thereby enhancing transparency for financial stakeholders. Such interpretability is pivotal for fostering trust in AI-driven fraud detection systems, as compliance teams and analysts can readily validate the rationale behind each flagged transaction.
IV-A Prospects for Future Studies